HIGH RESOLUTION FTIR EMISSIONE SPECTRA, MOLECULAR CONSTANTS, AND POTENTIAL CURVE OF GROUND-STATE GeO

\underline{EDWARD G. LEE}, JENNING Y. SETO, TSUYOSHI HIRAO, PETER F. BERNATH and ROBERT J. LE ROY, Guelph-Waterloo Centre for Graduate Work in Chemistry and Biochemistry, Department of Chemistry, University of Waterloo, Waterloo, Ontario N2L 3G1, Canada.

Extensive new high temperature, high resolution FTIR emission spectroscopy measurements for the five common isotopomers of GeO are combined with previous diode laser and microwave measurements in combined isotopomer analyses. New Dunham expansion parameters and an accurate analytical potential energy function are determined for the ground X^{1}Σ^{+} state.